The amplitude of two waves are in the ratio 5:2. if all other conditions for the two waves are same, then what is the ratio of their energy densities?
(a) 5 : 2
(b) 10 : 4
(c) 2.5 : 4
(d) 25 :4

Grade:12

2 Answers

Arun
3 years ago

As we learnt in 

Intensity of the wave -

I = \frac{1}{2}\rho \omega ^{2}A^{2}V

\rho = mass density

\omega = angular frequency

A = Amplitude

V = Wave speed

 

- wherein

flow of energy per unit area of cross section of the string in the unit time is known as intensity of wave.

 

 Energy density  \propto \:A^{2}

\Rightarrow \:\frac{E_{1}}{E_{2}}=(\frac{A_{1}}{A_{2}})^{2}=(\frac{5}{2})^{2}=25:4


 
Vikas TU
3 years ago

As sound wave travel through air, the element of air vibrate to produce change in density and pressure along the direction of motion of the wave.

- wherein

sound wave travel through any material medium with a speed that depend on the properties of medium.